OpenStudy (chelseatheruler):

How do you add sqrt(18) + sqrt(8)?

OpenStudy (therealmeeeee):

ok so sqrt 8 = sqrt 4 * sqrt 2 = 2*sqrt 2 sqrt 18 = sqrt 9 * sqrt 2 = 3*sqrt 2 2*sqrt2 + 3*sqrt2 = 5*sqrt2

OpenStudy (therealmeeeee):

Hope I help! :)

OpenStudy (***[isuru]***):

well .. u have to simply them a bit further first , u can write \[\sqrt{18 } = \sqrt{ 9 \times 2} = 3\sqrt{2}\] and \[\sqrt{8} = \sqrt{4 \times 2} = 2\sqrt{2}\] now u have something similer \[3\sqrt{2} + 2\sqrt{2}\] now to get the answer just add the numbers in front of the square root symbol which is \[5\sqrt{2}\]

OpenStudy (chelseatheruler):

Thanks again lol. :D

OpenStudy (***[isuru]***):

u r welcome :3